### v1

# Wide-Field, High-Resolution Reconstruction in Computational Multi-Aperture Miniscope Using a Fourier Neural Network

Wide-Field, High-Resolution Reconstruction in Computational Multi-Aperture Miniscope Using a Fourier Neural Network
[…]
journal: opticajournalarticletype: Research Article{abstract*}
Traditional fluorescence microscopy is constrained by inherent trade-offs among resolution, field-of-view, and system complexity. To navigate these challenges, we introduce a simple and low-cost computational multi-aperture miniature microscope, utilizing a microlens array for single-shot wide-field, high-resolution imaging. Addressing the challenges posed by extensive view multiplexing and non-local, shift-variant aberrations in this device, we present SV-FourierNet, a novel multi-channel Fourier neural network. SV-FourierNet facilitates high-resolution image reconstruction across the entire imaging field through its learned global receptive field. We establish a close relationship between the physical spatially-varying point-spread functions and the network’s learned effective receptive field. This ensures that SV-FourierNet has effectively encapsulated the spatially-varying aberrations in our system, and learned a physically meaningful function for image reconstruction. Training of SV-FourierNet is conducted entirely on a physics-based simulator. We showcase wide-field, high-resolution video reconstructions on colonies of freely moving C. elegans and imaging of a mouse brain section. Our computational multi-aperture miniature microscope, augmented with SV-FourierNet, represents a major advancement in computational microscopy and may find broad applications in biomedical research and other fields requiring compact microscopy solutions.
1 Introduction
Wide-field, high-resolution imaging plays a critical role across diverse scientific disciplines. Yet, traditional microscopy, which relies on single-objective lenses, is encumbered by a fundamental trade-off between field-of-view (FOV) and resolution[1]. To circumvent these constraints, array microscopes have been developed, facilitating rapid, wide-FOV image acquisition[2, 3, 4]. These systems utilize an ensemble of microscopes, each having its own lens and image sensor, to expand the FOV without sacrificing resolution. Despite these advancements, the scalability and miniaturization of array microscopes are limited by their dependence on bulky, specialized image sensor arrays, which pose significant challenges in applications demanding portability and compactness, such as endoscopy[5], on-chip microscopy[6], and in-vivo neural imaging [7].
To address these challenges, we introduce a computational multi-aperture miniature microscope (miniscope) that delivers wide FOV and micrometer resolution on a compact and lightweight device. Drawing inspiration from array microscopes, our design employs a microlens array (MLA) as the sole imaging component. Leveraging principles from lightfield and integral imaging techniques[8], our previous work has demonstrated single-shot 3D imaging capability using such a design[9, 10]. This work aims to provide a new perspective of this multi-aperture imaging system in order to address the challenges related to FOV and uneven resolution. Our approach is based on the premise that each lens within the MLA can capture high-quality image of a distinct, albeit limited, sub-FOV [Fig. 1(a)]. While individual lenses yield images that suffer from spatially-varying aberrations and contrast loss towards the edges of their respective sub-FOVs, combining information from various microlenses enables the reconstruction of a high-resolution image over an extended FOV, as illustrated in Fig. 1(a) and 1(c).
Figure 1: System overview. (a) Illustrations of the capability of individual lenses within the MLA, each focusing on a specific sub-FOV, with noticeable aberrations and diminishing contrast at the periphery. (b) The process of SV-FourierNet, where the captured image is first transformed into the Fourier domain and processed by multiple learnable filters for initial deconvolution, followed by information fusion and enhancement by a network to yield the final output. (c) The SV-FourierNet reconstruction achieves enhanced image quality and uniform resolution across an expanded FOV.
Departing from conventional array microscopes, our system utilizes a single image sensor to enhance device miniaturization and portability. Unlike other MLA-based multi-aperture imaging designs[11, 12, 13] that employ physical barriers to isolate sub-images captured by individual microlenses, our system opts for an unobstructed arrangement to maximize the effective imaging FOV. This design, however, introduces severe cross-talk between adjacent sub-images on the sensor, i.e. view multiplexing. Furthermore, the single-element MLA generates point spread functions (PSFs) that contain widely separated array foci and exhibit considerable spatial variability, i.e. non-local and spatially-varying PSFs. These complexities pose substantial challenges to achieving high-quality image reconstruction across the extended FOV.
Our solution to simultaneously tackle view multiplexing and non-local, spatially-varying aberrations is to employ a deconvolution algorithm. Traditional model-based methods rely on iterative algorithms [14, 15, 16], which are computationally intensive and time-consuming. Additionally, the reconstruction quality is greatly affected by manually tuned regularization terms.
Recent progress in deep-learning-based spatially-varying deconvolution methods have shown improvements in image quality and reconstruction speed[17, 18], emerging as a promising alternative to model-based methods. However, deep-learning methods often suffer from two major limitations in terms of scalability and locality bias. Specifically, the computational complexity and memory demands of convolutional neural networks (CNNs) increase with the input size. Thus, the direct application of CNNs to raw measurements containing tens of millions of pixels is impractical due to limited computational resources. The common workaround involves dividing raw images into smaller patches to make CNN training manageable. This method, however, assumes uniformity in the data distribution, and thus invariance in the PSFs for all image patches, overlooking the nuances due to spatially-varying aberrations. Consequently, such CNN-based methods often exhibit degraded resolution at FOV periphery, where aberrations diverge from those in central regions[10].
A second challenge in using CNNs for handling non-local, spatially-varying PSFs is the networks’ inherent locality bias [19, 20], a result of the small filter sizes in convolution layers that restrict CNN to capture global information. Although increasing CNN’s depth can broaden the “receptive field”, this approach escalates computational costs and complicates network training. The recently proposed MultiWienerNet [17] addresses these issues by initially applying deconvolution with multiple Wiener filters informed by a set of pre-calibrated, spatially-varying PSFs, followed by refinement through a CNN. Yet, this technique heavily depends on the choice of these initial PSFs that need to be sufficiently representative of the underlying spatially-varying aberrations across the FOV. An alternative approach is the Fourier Neural Network (FourierNet), which overcomes locality bias by applying filters in the Fourier domain [21, 22, 23, 24]. This strategy takes advantage of the global processing capabilities of the Fourier transform, allowing FourierNet to achieve a global receptive field with minimal layers, thus efficiently sidestepping the limitations faced by traditional CNNs. While FourierNet has been successfully applied in spatially invariant imaging systems [22, 23, 24], its ability to address spatially-varying systems remains unexplored.
In this work, we introduce Spatially-Varying FourierNet (SV-FourierNet) to facilitate wide-field, high-resolution reconstruction within our computational multi-aperture miniscope. We show that SV-FourierNet can be applied directly on full-FOV measurements containing 12-million pixels, circumventing the uneven resolution constraints typically associated with patch-based training methods. Furthermore, by employing multiple learnable Fourier filters [Fig. 1(b)], SV-FourierNet is effective in addressing the view multiplexing and highly non-local and spatially-varying array PSFs in our multi-aperture miniscope. The multi-channel Fourier layer in SV-FourierNet effectively performs a blind deconvolution, based on a low-rank imaging model, to process the input measurement. The resulting multi-channel deconvolved outputs are then fused and refined by an enhancement network, yielding the final reconstructed image.
To elucidate the working principle of SV-FourierNet, we employ a gradient-based method to probe the network’s “effective receptive field (ERF)”[25]. This method identifies the points in the measurement that most significantly affect the prediction of a specific point on the reconstructed object, closely resembling the PSF of our imaging system at the corresponding location. We show a close correspondence between SV-FourierNet’s ERFs and the physical PSFs at various field positions. This underscores that SV-FourierNet has encapsulated the spatially-varying aberrations in our system, and learned a physically meaningful function for image reconstruction.
Compared to current existing state-of-the-art deep learning methods[10, 17], SV-FourierNet demonstrates superior reconstruction quality, consistent resolution across the FOV, and enhanced inference speed. Trained entirely on a physics-based simulator, we show experimentally that SV-FourierNet attains a uniform resolution of 7.8 $\mu$m across a 6.5 mm FOV. Additionally, we present wide-field, high-resolution video reconstructions of freely moving C. elegans colonies and imaging of a weakly scattering mouse brain section.
Our computational multi-aperture miniature microscope augmented with SV-FourierNet represents a significant advancement in computational microscopy, promising wide-ranging applications in biomedical research and other areas requiring compact imaging solutions.
2 Methods
2.1 Setup of the multi-aperture miniscope
The multi-aperture miniscope is built with off-the-shelf and 3D-printed components. It uses an off-the-shelf plastic 3$\times$3 MLA (1 mm pitch, no. 630, Fresnel Technologies Inc.), forming a finite conjugate imaging geometry with a working distance of $\sim$7.5 mm and an MLA-sensor distance of $\sim$4.5 mm. For green fluorescence imaging, a hybrid filter set is incorporated, combining an interference filter (no. 535/50, Chroma Technology) placed in front of the MLA and a long-pass absorption filter (Wratten color filter no. 12, Edmund Optics) placed after the MLA. It uses a backside-illuminated (BSI) CMOS sensor (IM226, IDS Imaging) with 12 million pixels and 1.85 $\mu$m pixel size, achieving an effective pixel size of 3.7 $\mu$m. See more details in Supplement 1.
2.2 Spatially-varying imaging model
To bypass the need for extensive experimental data collection, we develop a highly efficient and accurate simulator to generate a comprehensive training dataset in silico. Building upon our previous work[10], this simulator incorporates a low-rank spatially-varying imaging model, improved with insights from a ray-tracing model, as detailed below.
We first collect a set of calibrated PSFs by scanning a single fluorescent bead across a 7 mm $\times$ 7 mm FOV with a uniform step size of 250 $\mu$m. To showcase the strong spatial variance of the PSFs, we calculate the Pearson Coefficient Correlation (PCC) map on the calibrated PSFs with respect to the on axis PSF, as shown in Fig. 2(a). The PCC map exhibits a rapid radial decay due to the limited angular response of the system and spatially-varying aberrations from the MLA. The central line profile shows that the PCC decreases to 0.2 with a 3.5 mm displacement.
Figure 2: Spatially-varying imaging model. (a) Spatial variance is quantified by the PCC map and the central line profile, which both exhibit a marked decline as the point source shifts away from the FOV center. (b) The low-rank model is computed on calibrated PSFs, which are cropped around each ML’s foci to enhance computational efficiency. The foci captured at different point source locations under the same ML show similar aberrations. TSVD is performed to compute the basis PSFs and coefficient maps for each ML. (c) Compared to the array decomposition method, the single-ML decomposition method reduces the model’s rank by 2$\times$ while maintaining accuracy. (d) The final imaging model involves computing $K$ weighted convolutions between the object and the basis PSFs for each ML and a superposition of all MLs to yield the multiplexed measurement. (e) The model is validated by comparing the simulated and physical measurements from the same object containing randomly distributed 10 $\mu$m fluorescent particles.
Next, we develop a low-rank model in three steps. To enhance computational efficiency, we first exploit the sparsity in the array PSF generated by the MLA and perform a cropping operation based on a ray tracing model to isolate each focal region (240 $\times$ 240 pixels) from the raw PSF measurement [Fig. 2(b)]. This method relies on the calibration of the chief ray of each microlens (ML) $(i,j)$ from an on-axis point source, which establishes the “anchor” coordinates in both the image space $(p_{x_{0}}^{i,j},q_{y_{0}}^{i,j})$ and the object space $(x_{0},y_{0})$. Given the depth $z$ and the point source’s lateral coordinates $(x,y)$, the corresponding PSF centroids for each ML $(i,j)$ $(p_{x}^{i,j},q_{y}^{i,j})$ is linearly proportional to the magnification $M_{z}$ and the object displacement:
$p_{x}^{i,j}=p_{x_{0}}^{i,j}-M_{z}\cdot(x-{x_{0}}),$ (1)
$q_{y}^{i,j}=q_{y_{0}}^{i,j}-M_{z}\cdot(y-{y_{0}}).$
Contrasting with cross-correlation methods typically employed to locate centroids of calibrated PSFs[10, 26], our ray tracing model embeds the image distortion information into the PSF patches. The distortion is characterized by the subsequent decomposition process and corrected during the reconstruction phase.
Second, to approximate the system’s spatial variance, we implement truncated singular value decomposition (TSVD) on the cropped foci for each individual ML, denoted as $H^{i,j}(p,q;x,y)$. Unlike our previous work[10] that performs a global decomposition on the array PSFs, our new method decomposes the foci of each ML independently. This is based on the observation that aberrations tend to be more consistent within a single ML but vary across MLs, largely due to the manufacturing variations, as shown in Fig. 2(b). Exploiting these distinct aberrations, our single-ML decomposition method effectively reduces the rank of the imaging model and computational complexity by 2×2$\times$2 × without compromising the model accuracy [Fig. 2(c)]. The decomposition process can be written as
$H^{i,j}(p,q;x,y)\approx\sum_{k=1}^{K}M^{i,j}_{k}(x,y)B^{i,j}_{k}(p,q),$ (2)
where $M^{i,j}_{k}(x,y)$ and $B^{i,j}_{k}(p,q)$ denote the $k^{\mathrm{th}}$ coefficient map and basis PSF for ML $(i,j)$, respectively. We select $K$ = 32, achieving an approximation error of $\sim$2.5on the calibration set.
Finally, the spatially-varying model is implemented in two steps, as shown in Fig. 2(d). First, the object is element-wise multiplied with a coefficient map and then convolved with the corresponding basis PSF for each ML. The $K$ weighted convolutions are summed to form a single-ML measurement. Next, all the single ML measurements are “placed back” according to the pre-calibrated anchor location $(p_{x_{0}}^{i,j},q_{y_{0}}^{i,j})$, and then summed together to construct the final multiplexed measurement. This process is described as
$g(p,q)=\sum_{i,j=1}^{3}\sum_{k=1}^{K}[M^{i,j}_{k}(p,q)\odot O(p,q)]\circledast_{p,q}B^{i,j}_{k}(p,q),$ (3)
where $O(p,q)$ is the object’s fluorescence distribution, and the image space coordinates $(p,q)$ are related to the object space coordinates $(x,y)$ by Eq. (1).
To expedite the simulation process, sub-images from individual MLs are computed in parallel, allowing the model to generate a measurement within 30 seconds. The sensor noise is sampled from a pre-calibrated mixed Poisson-Gaussian noise model and added as an augmentation to the measurements during the training process, which ensures that the synthetic measurements closely approximate real imaging conditions (See details in Supplement 1).
The spatially-varying imaging model is validated by comparing the simulated and experimental measurements on the same object containing randomly distributed particles, shown in Fig. 2(e) (See details in Supplement 1). The agreement between simulation and experiment, measured by PCC, achieves 0.92, indicating our model’s accuracy. Furthermore, two zoomed-in regions show that the model accurately reproduces the aberrations present in the physical measurement.
To promote generalization of SV-FourierNet, we simulate training data by applying our low-rank model to a diverse set of biological images collected from online repositories[27, 28, 29, 30, 31], including cells, vasculature, and brain sections, as well as synthetically generated fluorescent particles with random size, brightness, and locations, yielding around 5000 training images.
2.3 SV-FourierNet Design
SV-FourierNet features high scalability and effective processing of global spatial information, enabling the direct processing of raw measurements containing 12 million pixels. This capability allows it to simultaneously address the challenges of view multiplexing and non-local, spatially-varying aberrations in our system.
Figure 3: SV-FourierNet structure. The full measurement is first sent to the Fourier blind deconvolution module to perform initial reconstructions, followed by the enhancement module to form the final reconstruction.
SV-FourierNet comprises two core modules, including a blind Fourier deconvolution module and an enhancement module, as shown in Fig. 3. The measurement is first Fourier transformed into the frequency domain, followed by element-wise multiplication with multiple learnable complex Fourier filters, whose weights are initialized with a uniform distribution $U$[0, 1]. The filtered results are then inverse Fourier transformed back to the spatial domain, resulting in the initial multi-channel reconstructions. This blind deconvolution approach, in contrast to MultiWienerNet [17], bypasses the need for pre-selecting PSFs and facilitates the efficient learning of a low-rank Fourier basis. Additionally, by leveraging the conjugate symmetry of the Fourier transform, SV-FourierNet halves the number of training parameters required compared to those needed for learning directly in the spatial domain[17]. SV-FourierNet employs nine Fourier filters to collectively account for the spatial variance in our system.
The enhancement module employs a modified Residual Channel Attention Network (RCAN) [32], which includes a down-scaling module, a residual in residual (RIR) block and an up-scaling module. Each RIR block connects two residual groups with a long skip connection, facilitating information fusion across different scales. Channel attention modules embedded within the residual groups dynamically adjust the channel-wise feature weights, essential for synthesizing information from different Fourier channels. To prevent overfitting and to improve generalizability, the enhancement module is designed to be lightweight, with approximately 0.18 million parameters. Additional details about the network implementation is in Supplement 1.
The loss function is a sum of binary cross entropy (BCE) and Mean Squared Error (MSE):
$L_{\mathrm{total}}=L_{\mathrm{BCE}}+L_{\mathrm{MSE}},$ (4)
$L_{\mathrm{BCE}}=\frac{1}{n}\sum_{i=1}^{n}[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})],$ (5)
$L_{\mathrm{MSE}}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},$ (6)
where $y_{i}$ and $\hat{y}_{i}$ represent the true and predicted values, respectively, $i$ indexes the pixel and $n$ is the total number of pixels. This dual-loss approach leverages BCE for enhancing the reconstruction of sharp feature and MSE for capturing intensity variations, guiding the model towards accurate reconstructions in both structure and intensity. Gradient backpropagation from this loss updates the weights of both the Fourier filters and the enhancement module in an end-to-end manner.
We implement SV-FourierNet in PyTorch on an Nvidia A40 GPU. Training is conducted entirely on a simulated dataset with a batch size of 4. We preprocess the simulated measurement through cropping and padding operations to address the periodic ambiguity characteristic of array PSFs (see details in Supplement 1). Optimization employs the Adam optimizer alongside a cosine annealing scheduler, with the entire training process spanning approximately 48 hours.
2.4 Computation of the ERF
To assess SV-FourierNet’s ability to encapsulate non-local, spatially-varying aberrations and learn a physically meaningful inverse mapping, we calculate its ERF after the network is trained. This involves gradient backpropagation from a pixel on the reconstructed object to the input measurement[25], effectively measuring the influence of the measurement on a specific object pixel and generating a localized heatmap. In practice, this heatmap depends not only on the network’s trained weights, but also varies with different input objects, making it input-dependent. The ERF is the ensemble average of the heatmaps across a large dataset, effectively making it input-independent. The SV-FourierNet’s ERF is computed using a set of 520 test images.
We emphasize that the ERF-based method of visualizing the network’s response offers more insights into the network’s learned functional mapping compared to other techniques like saliency map[33] and Grad-CAM[34], which often produce more ambiguous heatmaps and are designed to operate on an individual input basis. The strength of our ERF-based analysis lies in its capacity to calculate a statistically averaged response, making the resulting heatmaps agnostic to the input. If the network fails to learn a meaningful mapping, the most likely ERF is a feature-less blur across a substantial portion of the input field, a consequence of the averaging process[25]. The highly localized ERFs suggests that SV-FourierNet has learned to generalize across different objects, and the learned mapping function aligns closely with the underlying imaging model.
3 Results
3.1 SV-FourierNet’s learned ERF aligns with physical PSFs
Figure 4: SV-FourierNet’s ERFs closely match the physically measured PSFs. The figures show the overlay between SV-FourierNet’s positive ERF components (in magenta) and the physical PSFs (in cyan) at various locations: (a) the FOV center, (b-c) near the FOV edges, and (d) beyond the training FOV. For the insets, the gray region represents a quarter of the circular FOV, marked by a grid with 1 mm intervals. The yellow marker denotes the specific position of the object point.
We observe that the ERFs of the network closely resemble the definition of the PSFs of the physical imaging system. To discern whether SV-FourierNet has learned a meaningful reconstruction function beyond mere image feature recognition, we directly compare SV-FourierNet’s ERFs with the system’s PSFs across different field locations.
SV-FourierNet’s ERFs contain both positive and negative components. The positive ERF components indicate that specific pixels in the measurement lead to an increase in intensity at the reconstructed pixel by SV-FourierNet. Given that PSFs in fluorescence imaging are inherently positive, it is the positive ERF components that parallel the physical PSFs’ function. As shown in Fig. 4, SV-FourierNet’s ERFs are highly localized and closely match the physically measured PSFs across the field. At the FOV center, the ERF aligns with the physical PSF, containing a widely distributed 3 $\times$ 3 foci array, as shown in Fig. 4(a). This demonstrates that SV-FourierNet has learned a global receptive field to synthesize information across the entire measurement and the network’s effective response agrees with the physical response of the imaging system. Near the FOV edges, due to the sensor’s finite size and the system’s limited angular response, the physical PSFs are truncated to 2 $\times$ 3 or 3 $\times$ 2 foci array respectively, as shown in Fig. 4(b) and 4(c). Nevertheless, the ERFs precisely localize the physical foci. This showcases SV-FourierNet’s capability to learn the severe spatial variance in the imaging model, including handling image truncation – a case that often induces numerical artifacts in model-based methods[35]. In addition, when assessing a point outside the SV-FourierNet’s trained FOV, the physical PSF appears as a truncated 2 $\times$ 2 foci array with noticeable contrast loss [Fig. 4(d)]. However, the ERF aligns precisely with this pattern, underscoring SV-FourierNet’s ability to learn a physically accurate inverse mapping. This ensures model generalization even beyond the trained FOV boundaries. Zoomed-in views of the foci regions further elucidate the severe aberrations and spatial variability of the PSFs, along with the precise co-localization between the ERFs and the physical PSFs.
The negative ERF components contribute to SV-FourierNet’s capability to demultiplex overlapping views in the measurement, essential for addressing the view-multiplexing challenge inherent to our system. These components, as detailed in Supplement 1, form a periodic array foci pattern corresponding to the MLA’s periodicity. This pattern arises from the measurement’s ambiguity caused by the MLA’s periodic structure, where displacement by a distance equal to the foci array’s period results in nearly identical measurements. The differences are only due to unique aberrations of the individual microlenses and the image truncation by the system. SV-FourierNet navigates this complexity by leveraging distinct aberration patterns across microlenses, accentuating the accurate object components through positive ERF components while suppressing contributions from ambiguous components via negative ERF components.
3.2 Demonstration of consistent resolution across a wide FOV
To showcase our system’s capability to achieve uniform resolution across a wide FOV, we characterize the resolution at both the FOV center and edges by imaging a fluorescence resolution target, as shown in Fig. 5. We also compare the reconstructions with the raw measurements from the central lens, which underscores the resolution enhancement and FOV expansion enabled by SV-FourierNet. In the central FOV, SV-FourierNet’s reconstruction surpass those from the central lens in both resolution and contrast, illustrating the network’s efficiency in effectively synthesizing information from multiple lenses for high-resolution reconstructions. At the FOV periphery, where the central lens exhibits pronounced aberration and vignetting artifacts, SV-FourierNet robustly reconstructs the target with high resolution, showcasing the network’s capability to account for strong spatially-varying aberrations. The line profiles further affirm SV-FourierNet’s capacity to achieve a consistent 7.8 $\mu$m resolution vertically and horizontally at both central and peripheral FOV regions.
Figure 5: Resolution characterization. SV-FourierNet reconstruction of fluorescent resolution target at the FOV center and edges, compared with measurements from the central lens to highlight the enhanced resolution and extended FOV achieved by our system. Line profiles confirms the system achieves a consistent 7.8 $\mu$m resolution both vertically and horizontally, at both FOV center and edges.
3.3 SV-FourierNet achieves state-of-the-art reconstruction quality and speed
[…]
Figure 6: Comparison of SV-FourierNet with existing state-of-the-art networks on simulated and experimental datasets. (a) Reconstruction results from simulated cell and vasculature measurements. SV-FourierNet achieves the best PSNR and SSIM for both types of samples. (b) Reconstruction results from experimental measurements on 10 $\mu$m fluorescent beads and freely moving C. elegans. SV-FourierNet provides the best visual quality and the fastest inference speed. Two C. elegans are indicated by the white and yellow arrows in both the measurement and the reconstructed images. The numbers on the measurement index the views from each microlens. SV-FourierNet successfully reconstructs the C. elegans populations across the entire FOV, while other networks suffer from loss of information especially near the FOV edges (marked by the white arrows).
To ensure a fair comparison, we employ the same loss function and training strategy across all models. Each model is trained on the same dataset for 48 hours. The models with the best SSIM are then chosen for the benchmark comparison. Moreover, the architecture of the enhancement network within MultiWienerNet is modified to be the same as that of SV-FourierNet. The detailed architecture of the benchmark networks can be found in Supplement 1.
[…]
MultiWienerNet improves reconstruction performance at the FOV edge through multiple initial Wiener deconvolutions. However, it fails to preserve high-resolution details, as shown in the zoomed-in areas. This underscores the advantage of learning inverse mapping of a spatially-varying model in the frequency domain over the spatial domain. The broad, smooth spectrum of the PSF, as opposed to the spatial domain’s sparse and sharp features, proves easier to learn and more resilient to potential errors possibly due to neural network’s spectral bias towards low-frequency features[36]. Frequency domain filters excel in addressing diverse frequency components, adeptly analyzing quasi-periodic signals from our MLA-based measurement and adapting to intricate noise characteristics. In contrast, Wiener deconvolution, optimized for uniform, additive white Gaussian noise, results in suboptimal initial reconstructions and the loss of fine details in complex scenarios.
[…]
Figure 7: Imaging of a freely-moving C. elegans colony. (a) SV-FourierNet’s reconstruction captures the rapid movement of the C. elegans across a 6.5 mm FOV. (b) The reconstructions agree well with the concurrent widefield measurements across the entire FOV and zoomed-in regions, demonstrating our system’s capacity to robustly image complex biological dynamics on a large scale.
3.4 Experimental demonstration on a colony of freely moving C. elegans
To showcase our system’s capability in imaging large-scale dynamic biological processes, we image colonies of freely moving C. elegans. Details about the sample preparation are provided in Supplement 1. The SV-FourierNet reconstruction [Fig. 7(a)] and concurrent widefield measurements [Fig. 7(b)] are compared.
We compare the full-FOV reconstruction of a single frame from SV-FourierNet with the corresponding widefield measurement. The reconstruction accurately recovers all the C. elegans, including those in central and edge FOV regions, showcasing consistent performance across the extensive FOV. To further demonstrate our system is robust to dynamic movements and varying local contrast, we compare sequences of SV-FourierNet reconstructions and widefield measurements for the area marked by the white box. These comparisons demonstrate that SV-FourierNet is capable of accurately recovering not only the larger, high-contrast C. elegans but also the younger specimens with faint fluorescence signals, as indicated by the yellow dashed box. Additionally, our system effectively recovers the dynamic movements of the C. elegans population, with a video shown in Visualization 1.
Figure 8: Imaging of a 75 $\mu$m thick brain section. (a) SV-FourierNet reconstruction (with the raw measurement as an inset) aligns with the widefield measurement from a 2$\times$ 0.1NA objective, showcasing the network’s ability for accurate reconstruction across an entire 6.5 mm brain section. (b) In the zoomed-in regions, SV-FourierNet reconstructions (with yellow image borders) agree with widefield measurements from a 10$\times$ 0.25NA objective (with red image borders), demonstrating our network’s capacity to achieve reconstruction at cellular resolution.
3.5 Experimental demonstration on fixed brain section
To demonstrate our system’s capability for high-quality imaging of complex biological tissues, we image a 75 $\mu$m thick, weakly scattering brain section containing GFP-expressing neurons in the bed nucleus of the stria terminalis (BNST). The BNST, crucial in regulating emotional and stress responses, is bilaterally located on either side of the brain, approximately 2.5 mm apart, with each region extending about 2 mm. These regions are characterized by their intricate spatial structure and varying neuron density levels (detailed sample information is in Supplement 1).
First, we evaluate the full-FOV SV-FourierNet reconstruction by comparing it with a 2$\times$ widefield measurement. Upon visual inspection, SV-FourierNet reliably reconstructs the bilateral structure of the labeled brain regions, as shown in Fig. 8(a). We then compare the reconstructions of four zoomed-in regions from both sides of the brain section with 10$\times$ widefield measurements (shown in Fig. 8(b)). SV-FourierNet accurately reconstructs neuronal structures from both densely and sparsely populated areas. These results demonstrate that SV-FourierNet, trained entirely on simulated data, can generalize to complex brain tissues, provide cellular-resolutions across the entire brain section, and is robust against weak tissue scattering.
4 Conclusion
We present a computational multi-aperture miniscope augmented with SV-FourierNet that performs single-shot high-resolution wide-FOV imaging. Our system is simple and compact, using only a single MLA for imaging. Unlike studies focusing on exploiting the 3D imaging capability with an MLA[9], we introduce a novel computational framework to significantly expand the FOV while ensuring uniform high resolution throughout. Our main innovation is SV-FourierNet, which employs multiple learnable Fourier filters to synthesize information over an extended FOV, addressing spatially-varying aberrations and view multiplexing inherent to our system. Furthermore, we demonstrate a novel utility of the network’s ERF in elucidating the mapping function it learns. We show that SV-FourierNet learns a physically meaningful reconstruction function for the spatially-varying imaging model, achieving uniform resolution throughout the extended FOV. Trained exclusively on simulated data, SV-FourierNet exhibits robust generalization to diverse experimental samples, as demonstrated on resolution targets, fluorescent beads, live C. elegans, and brain tissue sections. This underscores SV-FourierNet’s potential for a wide array of applications within biomedical research and beyond.
[…]
For the first time to our knowledge, we establish a connection between the ERF of a deconvolution network and the underlying physical imaging model. This enables a straightforward visualization of the function learned by the network, alleviating the “black box” aspect often associated with the deep learning networks. Furthermore, the ERF can offer a promising tool to inform network design, allow for predictions of network performance and alignment between the trained deconvolution model and the physical optical system by visualizing deviations of the ERFs from the physical PSFs.
Our accurate and efficient physics simulator eliminates the need for extensive and laborious physical data collection, facilitating a broad data distribution, which is particular impactful when ground truth is difficult to obtain. These simulated training data are crucial for the network’s ability to generalize robustly across different experimental data types. To broaden the platform’s application to more complex scattering biological tissues, it is possible to integrate the scattering process into the imaging model[38]. With training data derived from this enhanced imaging model, we anticipate the platform’s utility to more challenging biomedical imaging applications.
[…]
Funding This project has been made possible in part by National Institutes of Health (R01NS126596) and a grant from 5022 - Chan Zuckerberg Initiative DAF, an advised fund of Silicon Valley Community Foundation.
[…]
Acknowledgments The authors thank Boston University Shared Computing Cluster for proving the computational resources, Ian G. Davison, Kevin J Monk, and Brett T DiBenedictis for providing the brain slices, as well as Christopher A Gabel and Andrew S. Chang for providing the C. elegans sample. We also thank Boston University Computational Imaging Systems Lab group members for insightful discussions and Adelina Chau and Jamin Xie for their assistance with benchmark network tuning.
[…]
Disclosures The authors declare no conflicts of interest.
[…]
Data Availability The SV-FourierNet implementation, and the pretrained model are available at: https://github.com/bu-cisl/SV-FourierNet.
[…]
Supplemental document See Supplement 1 for supporting content.